OpenStudy (anonymous):

A house sold for $400,000. The mortgage on the house is $300,000. The documentary stamp tax on the deed is $0.70 per $100 or portion thereof. How much is the documentary stamp tax on the deed? $210 $280 $2,100 $2,800

OpenStudy (anonymous):

I got 2100 < @amistre64

OpenStudy (amistre64):

not really a biology question tho ...

OpenStudy (anonymous):

0.70/100 = 0.007 * 300,000 = 2100

OpenStudy (anonymous):

or 2800

OpenStudy (amistre64):

does the tax only apply to mortgage? or selling amount?

OpenStudy (anonymous):

I think the selling amount

OpenStudy (amistre64):

check your material, i think it relates to the mortgage

OpenStudy (amistre64):

"Documentary Stamp Tax at $.70 per $100.00, rounded up, based on the total purchase price" .. thats sorta vague to me

OpenStudy (amistre64):

i spose the total purchase price would be the down payment + loan .... so the selling price is apt

OpenStudy (amistre64):

400 000 100 -------- 4000 70 * 4000 100 7*400 = 2800

OpenStudy (anonymous):

ok thank you

OpenStudy (amistre64):

"doc stamp on the deed" is the key phrase ... they have differing doc stamps .... on the deed refers to total purchase price (amount paid for)